You have a firm that starts out with $60,000 in cash in the bank. You have three investment opportunities: (1) You can invest $30,000 today and get back a gross payoff of $45,000 next year; (2) you can invest $20,000 today and get back a payoff of $ 24,000 next year; (3) you can invest $10,000 today and get back a payoff of $20,000 next year. You can undertake any or all of these investment opportunities.

  1. Suppose the interest rate is 30%. What investment policy do you undertake, and what is the value of your firm today after you announce your investment policy?
  2. Suppose you want to consume $30,000 today and the rest next year. How much money do you need to borrow or lend? How much can you consume tomorrow?

Suppose a fourth investment opportunity (4) is added to the problem: i invest $15,000 today and get back $50,000 next year. In which projects should you invest?v
Answer #1

With 1st alternative we have return on investment is

=(45000/30000)-1=50%

2nd alternative return on investment

=(24000/20000)-1=20%

3rd alternative return on investment

=(20000/10000)-1=100%


Hence we should choose 3rd alternative because interest rate is 30% and we can earn highest returns by investing into 3rd alternative

Ans b)

IF we consume today $30000 by borrowing with 30% then we have to pay $39000 next year whenre we would invest $60000 into 3rd alternative then we have 60000(2)=120000

Hence we have $120000-$39000=$81000

Ans C)
if 4th alternative is available then we should invest into 4th alternative because (50000/15000)-1=333.33%

4th alternative is best